OpenStudy (ggray):

What is the value of 54-4a^2+3b^3 when a= -2 and b= 4? Will medal

OpenStudy (ggray):

Please?

OpenStudy (anonymous):

I believe it would be 230. Not sure. 54 - 4(-2^2) + 3(4^3). 54 - 4(4) + 3(64) = 54 - 16 + 192 246 - 16 = 230 This could be wrong! You can always ask someone else. xx

OpenStudy (ggray):

Thanks

OpenStudy (ggray):

@reneezrichter do you know?

OpenStudy (anonymous):

You're welcome, but I'm no professional, you may need to ask someone else in case they get different.

OpenStudy (ggray):

I'm pretty sure I did idk how to tag people to the question

OpenStudy (anonymous):

I got 230 :)

OpenStudy (ggray):

It was right but I didn't pass the test

OpenStudy (anonymous):

I'm sorry :(